Suppose a farmer in Georgia begins to grow peaches. He uses​ $1,000,000 in savings to purchase​ land, he rents equipment for ​$90,000 a​ year, and he pays workers ​$130,000 in wages. In​ return, he produces 200,000 baskets of peaches per​ year, which sell for ​$3.00 each. Suppose the interest rate on savings is 4 percent and that the farmer could otherwise have earned ​$45,000 as a shoe salesman.
What is the​ farmer's economic​ profit?

The peach farmer earns economic profit of ​$
nothing. ​(Enter your response as an​ integer.)

What is the​ farmer's accounting​ profit?
The peach farmer earns accounting profit of ​
Answer #1

Total revenue = 200000*3 = 600000

Economic profit = TR-(Explicit+Implicit costs)

= 600000-(90000+130000+45000+4% of 1000000)

= 600000-(305000)

= 295000

Accounting profit = TR-Explicit costs

= 600000-(90000+130000)

= 380000